OpenStudy (anonymous):

Christina keeps her DVD collection in a plastic tub. The table shows the number of each type of DVD in Christina's collection. Christina's DVD Collection Type of DVD Number of DVDs Sci-Fi 2 Action 3 Drama 8 Comedy 10 Romance 7 If two DVDs are chosen randomly one at a time, where the first DVD is replaced before the second DVD is chosen, what is the theoretical probability that both DVDs chosen are from the drama category?

OpenStudy (anonymous):

\[\left(\begin{matrix}8 \\ 2\end{matrix}\right)\left(\begin{matrix}20\\ 0\end{matrix}\right)/\left(\begin{matrix}28 \\ 2\end{matrix}\right)\]